# Partonic distribution functions and amplitudes using tensor network methods

### v2

@@ Abstract
Calculations of the parton distribution function (PDF) and distribution amplitude (DA) are highly relevant to core experimental programs as they provide non-perturbative inputs to inclusive and exclusive processes, respectively. Direct computation of the PDFs and DAs remains challenging because they are non-perturbative quantities defined as light-cone correlators of quark and gluon fields, and are therefore inherently time-dependent. In this work, we use a uniform quantum strategy based on tensor network simulation techniques to directly extract these hadronic quantities from first principles using the matrix product state of the hadrons in the same setup. We present exemplary numerical calculations with the Nambu-Jona-Lasinio model in 1+1 dimensions and compare with available exact diagonalization and quantum circuit simulation results. Using tensor networks, we evaluate the PDF and DA at various strong couplings in the large-qubit limit, which is consistent with expectations at perturbative and non-relativistic limits.
@@ 1 Introduction
Quantum computing technology has undergone rapid development and may provide an alternative path toward calculating the structure of hadrons from first principles Bauer:2022hpo ; DiMeglio:2023nsa . The exponential Hilbert space and quantum simulation algorithms provide a natural and efficient theoretical framework to simulate quantum field theory Jordan:2011ci ; Jordan:2014tma ; Jordan:2017lea . Studying real-time evolution. such as the case for the PDF and DA, is particularly suited to quantum simulation Lamm:2019uyc . The idea is to encode the fermionic and gluonic degrees of freedom into spin Hamiltonians and simulate the real-time matrix elements for evaluating these distribution functions and amplitudes directly on the quantum circuit. Recently, exploratory studies using quantum computing for parton structures are booming: from low-dimensional (1+1)-dimensional field theories Lamm:2019uyc ; Perez-Salinas:2020nem ; LiTianyin:2021kcs ; Li:2022lyt ; LiTianyin:2024nod ; Grieninger:2024cdl ; Grieninger:2024axp , worldline formulation Mueller:2019qqj , spacetime Wilson loop Echevarria:2020wct , to effective light-front Hamiltonian models Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp , these studies already exhibit interesting physics and potential. However, most quantum simulation for parton structure is still restricted in performing systematic calculation and limited in extending to the continuum, largely due to today’s noisy intermeidate-scale quantum infrastructure Preskill:2018jim .
@@
In this work, we aim to leverage tensor network (TN) methods in computing hadron structures. Traditionally used in condensed matter physics, TN have become an increasingly promising tool of its own to study hadronic physics and high-energy physics Silvi:2017srb ; Banuls:2018jag . Owing to its similarity with quantum computing, TN, despite being a classical computation approach, is used as a natural testbed for quantum simulation, especially for low dimensional and low entanglement problems. In particular, TN is an efficient simulation strategy to simulate quantum-many body systems by representing the quantum states locally in the Hilbert space and their dynamics can often be bounded by an area-law of entanglement entropy Orus:2013kga ; mps-reps-gs . Importantly, TN is free from sign problems and reliable in accessing regimes inaccessible to Monte Carlo methods Banuls:2018jag . Recently, TN methods are successfully demonstrated in evaluations of hadronic vacuum polarization Barata:2024bzk and the PDF Banuls:2024oxa in the 1+1 QED for the first time. Here, we aim to compute both the PDF and DA in the same setting of the Nambu-Jona-Lasinio (NJL) model in 1+1 dimensions using efficient TN simulation algorithms. Having direct access to hundreds of qubits with TN, we present numerical simulation at large-qubit limit that is representative of the continuum limit. Specifically, we compare the behavior of the PDF and DA at various mass values and strong coupling strengths to make qualitative comparison with available results in perturbative QCD and the non-relativistic limits.
@@ 2.1 Parton distribution functions and amplitudes
$f_{q/h}(x)=$ $\int_{-\infty}^{\infty}\frac{\mathrm{d}t}{4\pi}e^{-itx\vec{n}\cdot\vec{P}}\bra{h(\vec{P})}\bar{\psi}(t\vec{n})W(t\vec{n}\leftarrow\vec{0})\vec{n}\cdot\vec{\gamma}\psi(\vec{0})\ket{h(\vec{P})}\,,$ (1)
@@
$\phi_{h}(x)=$ $\frac{1}{f_{h}}\int_{-\infty}^{\infty}\mathrm{d}t\,\,e^{-it(x-1)\vec{n}\cdot\vec{P}}\bra{\Omega}\bar{\psi}(t\vec{n})W(t\vec{n}\leftarrow\vec{0})\vec{n}\cdot\vec{\gamma}\psi(\vec{0})\ket{h(\vec{P})}\,,$ (3)
@@
2.2 Nambu-Jona-Lasinio model in 1+1 dimensions
Simulating the 3+1 QCD Hamiltonian for hadronic structures is presently very challenging; therefore in this work we focus on the Nambu-Jona-Lasinio (NJL) model in 1+1 dimensions to demonstrate the computation of the PDF and DA in a common footing. The same procedure can also be applied to hadronic tensors. The NJL model is an effective model for low-energy two-flavored QCD Nambu:1961tp ; Nambu:1961fr and its simplified and renormalizable version, the Gross-Neveu (GN) model Gross1974 , is an ideal candidate for illustrative calculations including chiral symmetry breaking Shi:2015ufa , chiral condensate Huang:2001yw , phase transitions Thies:2019ejd ; Czajka:2021yll , and so forth. We start with the Lagrangian density of the GN model with zero chemical potential, given by,
@@ 2.2 Nambu-Jona-Lasinio model in 1+1 dimensions
$\begin{split}H&=\frac{1}{2a}\left[\sum_{n=0}^{N-2}\left(-\sigma^{+}_{n}\sigma^{z}_{n}\sigma^{-}_{n+1}+\sigma^{+}_{n+1}\sigma^{z}_{n}\sigma^{-}_{n}\right)\right]+m\sum_{n=0}^{N-1}(-1)^{n}\sigma^{+}_{n}\sigma^{-}_{n}\\ &-\frac{g}{a}\sum_{n=0}^{N/2-1}\left(\sigma^{+}_{2n}\sigma^{-}_{2n}-\sigma^{+}_{2n+1}\sigma^{-}_{2n+1}\right)^{2}\,,\end{split}$ (13)
@@
Now we compute the PDF in eq. (1) and DA in eq. (3) using the obtained (1+1)-dimensional NJL Hamiltonian. Specifically, we use $\gamma^{+}=\frac{1}{\sqrt{2}}(\gamma^{0}+\gamma^{1})=\frac{1}{\sqrt{2}}(\sigma^{z}-i\sigma^{y})$ and $\vec{n}=(1,-1)$ for the light-cone coordinate. It follows that $\vec{n}\cdot{\vec{\gamma}}=\sqrt{2}\gamma^{+}$ and $\vec{n}t=(1,-1)z$ with a reparameterized variable $z$ for space and time. Since the gauge fields are absent in the NJL model, the explicit Wilson lines in the PDF and DA vanish and we are left with time evolution operators of the Hamiltonians. While this feature of the NJL model dramatically simplifies the calculation, we are well aware that one should include dynamical gluon contributions in a more realistic scenario with available methods for simulating the Wilson line Zohar:2012ts ; Brennen:2015pgn ; Lamm:2018siq ; Zohar:2019tsa ; Echevarria:2020wct , which we leave for a subsequent investigation. Recently, treatment of evolution of the Wilson lines was proposed and detailed for the Schwinger model Banuls:2024oxa . Applying the same JW encoding to field operators $\psi^{\dagger}$ and $\psi$, the final expression of the PDF and DA are related to the sum of four matrix element expectations,
$f_{q/h}(x)=$ $\frac{1}{4\pi}\sum_{z=\bar{N}_{\mathrm{min}}}^{\bar{N}_{\mathrm{max}}}e^{-ixM_{h}z}\bra{h}\big{(}M_{00}(z)+M_{11}(z)-M_{01}(z)-M_{10}(z)\big{)}\ket{h}\,,$ (14)
$\phi_{h}(x)=$ $\frac{1}{f_{h}}\sum_{z=\bar{N}_{\mathrm{min}}}^{\bar{N}_{\mathrm{max}}}e^{-i(x-1)M_{h}z}\bra{\Omega}\big{(}M_{00}(z)+M_{11}(z)-M_{01}(z)-M_{10}(z)\big{)}\ket{h}\,,$ (15)
@@
$M_{ij}(z)$ $=e^{iHz}\chi^{\dagger}_{-2z+i+\bar{N}}e^{-iHz}\chi_{j+\bar{N}}\,,$ (16)
with evolution time $z=na/2$ relating to sites $n$ and $n+1$, and $\bar{N}_{\mathrm{min}}$ and $\bar{N}_{\mathrm{max}}$ mark the boundary of the lattice. Here, we always center the system around $x=0$ positions for the simulation, so $\bar{N}_{\mathrm{min}}=-\lfloor\bar{N}/2\rfloor$ and $\bar{N}_{\mathrm{max}}=-\lfloor\bar{N}/2\rfloor+\bar{N}-1$ with the center qubit index $\bar{N}$. Then, for a sufficiently large $\bar{N}$, our setup is a representation of the physical system. In practice, when calculating the PDF, we also subtract the vacuum expectation from the hadron expectation due to finite system size used in our simulation Collins:2011zzd .
@@ 2.3 Tensor network methods
Tensor networks provide a convenient alternative for many-body simulations, especially since today’s quantum computers are still limited both in the number and fidelity of available qubits. TN methods recast (either exactly or up to accurate approximation) many-body quantum systems as a matrix product state (MPS) mps-reps-gs and use various algorithms and variational methods for studying the dynamics of MPS systems. MPS’s are especially well suited for 1D spatial systems as entropy area laws force an upper bound on entanglement vidal-erg . For this reason, TN methods provide efficient many body quantum simulations with classical hardware, especially for calculating hadronic structures in the 1+1 dimensional model considered in this work [^2] Note, for the same reason, TN do not replace standard QC methods but complement them..
@@ 2.3.1 Density matrix renormalization group
Preparation of the ground state or excited states is often the starting point for any quantum simulation Jordan:2011ci . While many quantum simulation methods have been suggested, including adiabatic quantum computation Chakraborty:2020uhf , Hamiltonian variational algorithm wiersema2020exploring , quantum subspace expansion mcclean2017hybrid , Rodeo algorithm Choi:2020pdg , among others; it is still typically very complicated to prepare the states on quantum circuits. In TN, by comparison, this complication is largely reduced utilizing efficient representation of the Hilbert space using MPS. An MPS is a many-body wave function decomposed into local tensors via singular value decomposition. In this work, we use the density matrix renormalization group (DMRG) algorithm to obtain the ground state and first excited state in MPS form White:1992zz ; White:1993zza ; Schollwock2005 ; schollwock2011density ; frank-dmrg .
The DMRG method basically takes the ansatz MPS and calculates the lattice configuration with the least amount of energy. It starts with an initial wave function, and then the energy is minimized by freely moving around two lattice sites while fixing all others. This process is repeated for many new lattice sites until convergence occurs. Traditionally, DMRG is used to study ground states of one-/two-dimensional many body systems frank-dmrg . The computational time scaling of the DMRG is $\mathcal{O}(N\chi^{3})$, where $\chi$ is bond dimension and $N$ is system size, i.e., total number of qubits dmrg-on-tpul . Specifically, the bond dimension is the dimension of the connecting index between two local tensors of an MPS. For calculating the first excited state, we construct a modified Hamiltonian that consists of the original Hamiltonian and a projection of the previously calculated ground state wave function. The projection has a coefficient “penalty” term that penalizes overlaps between the excited and ground states, which ensures orthogonality and minimization of the DMRG outside of the ground state ITensor:2022 . In this way, we use the DMRG algorithm to obtain both the ground state and the first excited state as MPS for the NJL model.
@@ 2.3.2 Time dependent variational principle
To calculate the matrix elements for the PDF and DA, as in eq. (16), we evolve the prepared MPS state in real time and collapse with a final MPS. The two primary algorithms for the time evolution of the MPS are the time-evolving block decimation (TEBD) Vidal:2003lvx ; verstraete2004matrix and time dependent variational principle (TDVP) Haegeman:2011zz ; haegeman2016unifying . Both of them have the same computational time scaling of $\mathcal{O}(N\chi^{3})$ as the DMRG algorithm, where $\chi$ is the bond dimension and $N$ is the system size. For time step $\tau$, the TDVP and TEBD algorithms have $O(\tau^{2})$ error scaling for the lower order dmrg-on-tpul ; tdvp-scaling ; vidal-tebd-many-body .
In principle, the TEBD method breaks up the Hamiltonian into small chunks and applies imaginary time evolution via Trotter gates that account for near neighbor interactions. The entanglement grows exponentially with time and TEBD method loses accuracy for long time evolution vidal-tebd-qc ; vidal-tebd-many-body . Specifically, for an $N$-qubit system, TEBD algorithm requires $\mathcal{O}(N\exp{(E_{\chi})})$ parameters where $E_{\chi}$ is the entanglement entropy, that is defined as von Neumann entropy $E_{\chi}=-\operatorname{Tr}(\rho\log(\rho))$ of the reduced density matrix $\rho$ from the whole quantum system to a block of local sites vidal-erg . If the bond dimension is not allowed to grow to account for this, then truncation errors occur vidal-tebd-qc ; vidal-tebd-many-body .
On the other hand, the TDVP algorithm projects the MPS onto a tangent space manifold, on which it solves the time dependent Schrodinger equation. This method doesn’t have the entanglement and error growth that Trotter gates have for larger systems, as it does not use Trotter gates at all. Instead, TDVP solves the Schrodinger equation projected onto a manifold, often chosen via the global Krylov method, which globally maintains all interaction terms Haegeman:2011zz . Potential errors are not as controllable, but are reduced with the use of Krylov time evolution itensor-tdvp . Though the TDVP and TEBD algorithms have the same scaling up to a prefactor, the TDVP algorithm has both a smaller prefactor for computational time scaling and less bond dimension growth Haegeman:2011zz ; itensor-tdvp . In this work, we primarily used TDVP, as it showed better performance than TEBD for NJL model, espeicially for high qubit simulations.

### v3

@@ Abstract
Calculations of the parton distribution function (PDF) and distribution amplitude (DA) are highly relevant to core experimental programs as they provide non-perturbative inputs to inclusive and exclusive processes, respectively. Direct computation of the PDFs and DAs remains challenging because they are non-perturbative quantities defined as light-cone correlators of quark and gluon fields, and are therefore inherently time-dependent. In this work, we use a uniform quantum strategy based on tensor network simulation techniques to directly extract these hadronic quantities from first principles using the matrix product state of the hadrons in the same setup. We present exemplary numerical calculations with the Nambu–Jona-Lasinio model in 1+1 dimensions and compare with available exact diagonalization and quantum circuit simulation results. Using tensor networks, we evaluate the PDF and DA at various strong couplings in the large-qubit limit, which is consistent with expectations at perturbative and non-relativistic limits.
@@ 1 Introduction
Quantum computing technology has undergone rapid development and may provide an alternative path toward calculating the structure of hadrons from first principles Bauer:2022hpo ; DiMeglio:2023nsa . The exponential Hilbert space and quantum simulation algorithms provide a natural and efficient theoretical framework to simulate quantum field theory Jordan:2011ci ; Jordan:2014tma ; Jordan:2017lea . Studying real-time evolution, such as the case for the PDF and DA, is particularly suited to quantum simulation Lamm:2019uyc . The idea is to encode the fermionic and gluonic degrees of freedom into spin Hamiltonians and simulate the real-time matrix elements for evaluating these distribution functions and amplitudes directly on the quantum circuit. Recently, exploratory studies using quantum computing for parton structures are booming: from low-dimensional (1+1)-dimensional field theories Lamm:2019uyc ; Perez-Salinas:2020nem ; LiTianyin:2021kcs ; Li:2022lyt ; LiTianyin:2024nod ; Grieninger:2024cdl ; Grieninger:2024axp , worldline formulation Mueller:2019qqj , spacetime Wilson loop Echevarria:2020wct , to effective light-front Hamiltonian models Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp , these studies already exhibit interesting physics and potential. However, most quantum simulation for parton structure is still restricted in performing systematic calculation and limited in extending to the continuum, largely due to today’s noisy intermediate-scale quantum infrastructure Preskill:2018jim .
@@
In this work, we aim to leverage tensor network (TN) methods in computing hadron structures. Traditionally used in condensed matter physics, TN have become an increasingly promising tool of its own to study hadronic physics and high-energy physics Silvi:2017srb ; Banuls:2018jag . Owing to its similarity with quantum computing, TN, despite being a classical computation approach, is used as a natural testbed for quantum simulation, especially for low dimensional and low entanglement problems. In particular, TN is an efficient simulation strategy to simulate quantum-many body systems by representing the quantum states locally in the Hilbert space and their dynamics can often be bounded by an area-law of entanglement entropy Orus:2013kga ; mps-reps-gs . Importantly, TN is free from sign problems and reliable in accessing regimes inaccessible to Monte Carlo methods Banuls:2018jag . Recently, TN methods are successfully demonstrated in evaluations of hadronic vacuum polarization Barata:2024bzk and the PDF Banuls:2024oxa ; Banuls:2025wiq in the 1+1 QED for the first time. Here, we aim to compute both the PDF and DA in the same setting of the Nambu–Jona-Lasinio (NJL) model in 1+1 dimensions using efficient TN simulation algorithms. Having direct access to hundreds of qubits with TN, we present numerical simulation at large-qubit limit that is representative of the continuum limit. Specifically, we compare the behavior of the PDF and DA at various mass values and strong coupling strengths to make qualitative comparison with available results in perturbative QCD and the non-relativistic limits.
@@ 2.1 Parton distribution functions and amplitudes
[…]
@@
[…]
@@
2.2 Nambu–Jona-Lasinio model in 1+1 dimensions
Simulating the (3+1)-d QCD Hamiltonian for hadronic structures is presently very challenging; therefore in this work we focus on the Nambu–Jona-Lasinio (NJL) model in 1+1 dimensions to demonstrate the computation of the PDF and DA in a common footing. The same procedure can also be applied to hadronic tensors. The NJL model is an effective model for low-energy two-flavored QCD Nambu:1961tp ; Nambu:1961fr and its simplified and renormalizable version, the Gross-Neveu (GN) model Gross1974 , is an ideal candidate for illustrative calculations including chiral symmetry breaking Shi:2015ufa , chiral condensate Huang:2001yw , phase transitions Thies:2019ejd ; Czajka:2021yll , and so forth. We start with the Lagrangian density of the GN model with zero chemical potential, given by,
@@ 2.2 Nambu–Jona-Lasinio model in 1+1 dimensions
[…]
@@
Now we compute the PDF in eq. (1) and DA in eq. (3) using the obtained (1+1)-dimensional NJL Hamiltonian. Specifically, we use $\gamma^{\pm}=\frac{1}{\sqrt{2}}(\gamma^{0}\pm\gamma^{1})=\frac{1}{\sqrt{2}}(\sigma^{z}\pm i\sigma^{y})$ and $\vec{n}=(1,1)$ for the light-cone coordinate. It follows that $\vec{n}\cdot{\vec{\gamma}}=\sqrt{2}\gamma^{-}$ and $\vec{n}t=(1,1)z$ with a reparameterized variable $z$ for space and time. Since the gauge fields are absent in the NJL model, the explicit Wilson lines in the PDF and DA vanish and we are left with time evolution operators of the Hamiltonians. While this feature of the NJL model dramatically simplifies the calculation, we are well aware that one should include dynamical gluon contributions in a more realistic scenario with available methods for simulating the Wilson line Zohar:2012ts ; Brennen:2015pgn ; Lamm:2018siq ; Zohar:2019tsa ; Echevarria:2020wct , which we leave for a subsequent investigation. Recently, treatment of evolution of the Wilson lines was proposed and detailed for the Schwinger model Banuls:2025wiq . Applying the same JW encoding to field operators $\psi^{\dagger}$ and $\psi$, the final expression of the PDF and DA are related to the sum of four matrix element expectations,
[…]
@@
$M_{ij}(z)$ $=e^{iHz}\chi^{\dagger}_{2z+i+\bar{N}}e^{-iHz}\chi_{j+\bar{N}}\,$, (16)
with evolution time $z=na/2$ relating to sites $n$ and $n+1$, and $\bar{N}_{\mathrm{min}}$ and $\bar{N}_{\mathrm{max}}$ mark the boundary of the lattice. Here, we always center the system around $x=0$ positions for the simulation, so $\bar{N}_{\mathrm{min}}=-\lfloor\bar{N}/2\rfloor$ and $\bar{N}_{\mathrm{max}}=\lfloor\bar{N}/2\rfloor$. We then have that qubits $\bar{N},\bar{N}+1$ simulate $x=0$. Then, for a sufficiently large $\bar{N}$, our setup is a representation of the physical system. In practice, when calculating the PDF, we also subtract the vacuum expectation from the hadron expectation due to finite system size used in our simulation Collins:2011zzd .
@@ 2.3 Tensor network methods
Tensor networks provide a convenient alternative for many-body simulations, especially since today’s quantum computers are still limited both in the number and fidelity of available qubits. TN methods recast (either exactly or up to accurate approximation) many-body quantum systems as a matrix product state (MPS) mps-reps-gs and use various algorithms and variational methods for studying the dynamics of MPS systems. MPS is particularly well-suited for representing the ground states of local, gapped Hamiltonians in 1D systems with a unique ground state and finite interaction strengths, where entropy area laws impose an upper bound on entanglement Hastings:2007iok ; vidal-erg . For this reason, TN methods provide efficient many body quantum simulations with classical hardware for calculating hadronic structures in the 1+1 dimensional model considered in this work [^2] Note, for the same reason, TN do not replace standard QC methods but complement them..
@@ 2.3.1 Density matrix renormalization group
Preparation of the ground state or excited states is often the starting point for any quantum simulation Jordan:2011ci . While many quantum simulation methods have been suggested, including adiabatic quantum computation Chakraborty:2020uhf , Hamiltonian variational algorithm wiersema2020exploring , quantum subspace expansion mcclean2017hybrid , Rodeo algorithm Choi:2020pdg , among others; it is still typically very complicated to prepare the states on quantum circuits. In TN, by comparison, this complication is largely reduced utilizing efficient representation of the Hilbert space using MPS. An MPS is an ansatz for the many-body wave function composed of local tensors,
[…]
Each $A_{i}^{s_{i}}$ is a matrix associated with site $i$ and state $s_{i}$ with size $\chi\times\chi$ where $\chi$ is the bond dimension between site $i$ and $i+1$. For open boundary conditions, $A_{1}^{s_{1}}$ is a $1\times\chi$ row vector and $A_{N}^{s_{N}}$ is a $\chi\times 1$ column vector. The bond dimension $\chi$ determines the entanglement capacity; a larger $\chi$ captures more correlations but increases computational cost.
[…]
Traditionally, the variational algorithm is used to study ground states of one-/two-dimensional many body systems frank-dmrg . The computational time scaling is $\mathcal{O}(N\chi^{3})$, where $\chi$ is bond dimension and $N$ is system size, i.e., total number of qubits dmrg-on-tpul . Specifically, the bond dimension is the dimension of the connecting index between two local tensors of an MPS. For calculating the first excited state, we construct a modified Hamiltonian that consists of the original Hamiltonian and a projection of the previously calculated ground state wave function. The projection has a coefficient “penalty” term that penalizes overlaps between the excited and ground states, which ensures orthogonality and minimization outside of the ground state ITensor:2022 . In this way, we use the variational algorithm to obtain both the ground state and the first excited state as MPS for the NJL model.
@@ 2.3.2 Time dependent variational principle
To calculate the matrix elements for the PDF and DA, as in eq. (16), we evolve the prepared MPS state in real time and collapse with a final MPS. The two primary algorithms for the time evolution of the MPS are the time-evolving block decimation (TEBD) Vidal:2003lvx ; verstraete2004matrix and time dependent variational principle (TDVP) Haegeman:2011zz ; haegeman2016unifying . Both of them have the same computational time scaling of $\mathcal{O}(N\chi^{3})$ as the variational search algorithm, where $\chi$ is the bond dimension and $N$ is the system size. For time step $\tau$, the TDVP and TEBD algorithms have $O(\tau^{2})$ error scaling for the lower order dmrg-on-tpul ; tdvp-scaling ; vidal-tebd-many-body .
In principle, the TEBD method breaks up the Hamiltonian into small chunks and applies real or imaginary time evolution via Trotter gates that account for near neighbor interactions vidal-tebd-qc ; vidal-tebd-many-body . The entanglement, in the worst case, can grow linear in time, which means the bond dimension of the MPS has to grow exponentially over time for a faithful representation of the state Kim:2013etb ; Schuch_2008 ; Calabrese:2005in ; Schachenmayer:2013 . In general, the entanglement growth is highly problem dependent. For gapped systems that evolve adiabatically close to the ground state at any time, typically it does not lead to a significant growth of entanglement, while out-of-equilibrium dynamics can trigger the worst case scenario Orus:2013kga . If the bond dimension is not allowed to grow to account for this, then truncation errors occur vidal-tebd-qc ; vidal-tebd-many-body .
On the other hand, the TDVP algorithm projects the MPS onto a tangent space manifold, on which it solves the time dependent Schrodinger equation. TDVP solves the Schrodinger equation projected onto a manifold, often chosen via the global Krylov method, which globally maintains all interaction terms Haegeman:2011zz . Potential errors are not as controllable, but are reduced with the use of Krylov time evolution itensor-tdvp . Though the TDVP and TEBD algorithms have the same scaling up to a prefactor, the TDVP algorithm has both a smaller prefactor for computational time scaling and less bond dimension growth Haegeman:2011zz ; itensor-tdvp . In this work, we primarily used TDVP, as it showed better performance than TEBD for NJL model, espeicially for high qubit simulations.